OpenStudy (anonymous):

a television game show has 4 doors of which contestant must pick two. behind two of the doors are expensive cars and behind the other 2 doors are consolation prizes. find the probability that the contestant wins both cars, no cars, or at least 1 car

OpenStudy (anonymous):

both cars \[\frac{2}{4}\times \frac{1}{3}\]

OpenStudy (anonymous):

no cars is the same \[\frac{1}{6}\]

OpenStudy (anonymous):

at least one car is \[1-\frac{1}{6}=\frac{5}{6}\]